OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3).

OpenStudy (anonymous):

OK, step one, what's the slope of the line we're trying to be parallel to?

OpenStudy (anonymous):

That's all the question gives me, if that's what you're asking..

OpenStudy (anonymous):

No, I'm saying, can you inspect the equation y=7x+2 and tell me the slope of that line?

OpenStudy (anonymous):

The answer, I'm totally lost. lol

OpenStudy (anonymous):

ok the slope intercept formula is y=mx+b where m is the slope and b is (0,b0 the y-intercept

OpenStudy (anonymous):

sorry where b is (0,b) the y-intercept

OpenStudy (anonymous):

y = 7x - 10

OpenStudy (anonymous):

if they give u an equation and they ask you find a parallel line to it ... you will use the slope the same slope the gave you that means m1=m2

OpenStudy (anonymous):

and will use the points to find b in the equation... lets see... y=7x+b if we use the points to find b we will get (-3)=7(1)+b -3=7+b -7 -7 -10=b

OpenStudy (anonymous):

so your equation for a parallel line wil be y=7x-10